# Alpamayo-R1: Bridging Reasoning and Action Prediction for Generalizable Autonomous Driving in the Long Tail

### v1

@@ Abstract
End-to-end architectures trained via imitation learning have advanced autonomous driving by scaling model size and data, yet performance remains brittle in safety-critical long-tail scenarios where supervision is sparse and causal understanding is limited. To address this, we introduce Alpamayo-R1 (AR1), a vision–language–action model (VLA) that integrates Chain of Causation reasoning with trajectory planning to enhance decision-making in complex driving scenarios. Our approach features three key innovations: (1) the Chain of Causation (CoC) dataset, built through a hybrid auto-labeling and human-in-the-loop pipeline producing decision-grounded, causally linked reasoning traces aligned with driving behaviors; (2) a modular VLA architecture combining Cosmos-Reason, a Vision-Language Model pre-trained for Physical AI applications, with a diffusion-based trajectory decoder that generates dynamically feasible plans in real time; (3) a multi-stage training strategy using supervised fine-tuning to elicit reasoning and reinforcement learning (RL) to optimize reasoning quality via large reasoning model feedback and enforce reasoning-action consistency. Evaluation shows AR1 achieves up to a 12% improvement in planning accuracy on challenging cases compared to a trajectory-only baseline, with a 35% reduction in off-road rate and 25% reduction in close encounter rate in closed-loop simulation. RL post-training improves reasoning quality by 45% as measured by a large reasoning model critic and reasoning=action consistency by 37%. Model scaling from 0.5B to 7B parameters shows consistent improvements. On-vehicle road tests confirm real-time performance (99 ms latency) and successful urban deployment. By bridging interpretable reasoning with precise control, AR1 demonstrates a practical path towards Level 4 autonomous driving. We plan to release AR1 models and a subset of the CoC in a future update.
@@ 1 Introduction
The evolution of autonomous driving systems has witnessed a paradigm shift from traditional modular architectures (Urmson et al., 2008; Paden et al., 2016; Zhang et al., 2018; Lefèvre et al., 2014) to end-to-end (E2E) driving frameworks (Bojarski et al., 2016; Hu et al., 2023; Gu et al., 2023; Weng et al., 2024; Wu, 2025), a transition increasingly embraced by industry. In contrast to modular designs that explicitly separate perception, prediction, and planning with hand-crafted intermediate representations, E2E approaches map raw sensor inputs directly to vehicle motion through jointly trained neural networks. This unified formulation eliminates manually engineered interfaces, enabling joint optimization and data-driven policy learning at scale. Recent advances in transformer-based architectures, coupled with large-scale driving datasets have further improved the overall performance and generalization of the E2E driving paradigm. Despite these successes, current E2E approaches remain fragile in handling long-tail and safety-critical situations, where sparse supervision and the need for high-level reasoning pose significant challenges.. Consequently, a significant gap persists between the capabilities of existing E2E models and the requirements for achieving robust Level-4 autonomy with driving-specific reasoning capabilities.
@@ 3 Building a Reasoning VLA Architecture
with each component conditioned on all previous ones. By default, the model is trained to predict the entire 6s-long future trajectory sequence
@@ 3.2.1 Vision Encoding
As an example, for $S_{x}=S_{y}=96,S_{z}=48$, and $p_{x}=p_{y}=p_{z}=8$, only 288 tokens are needed to represent one timestep of observations, irrespective of the number of cameras or their resolution. For a 7-camera vehicle setup, this equates to approximately $41.1$ tokens per image ($3.9\times$ less than single-image tokenization). Further, as we will show in Sec.˜6.6, this efficiency is achieved without major compromises to end-to-end driving metrics.
@@
Accordingly, AR1 also supports using multi-camera video tokenizers that directly encode entire sequences of camera observations from multiple timesteps. One example is Flex (Yang et al., 2025), which compresses a set of image tokens from multiple cameras and timesteps via full self-attention layers and a fixed set of query vectors, providing an explicit mechanism to control the magnitude of the information bottleneck. As will be shown in Sec.˜6.6, this approach can achieve an up to $20\times$ token compression rate (compared to single-image tokenization) while maintaining or even improving downstream driving metrics.
@@ 5.1 Action Modality Injection
Similar to $\pi_{0.5}$-KI (Driess et al., 2025), we adopt a separate action-expert to decode actions via flow matching (Janner et al., 2022; Lipman et al., 2023; Zhong et al., 2023; Jiang et al., 2023). The action-expert follows the same Transformer architecture as the VLM, using the same number of attention heads and attention dimensions, but with a smaller hidden embedding size and MLP dimension for efficiency. At each diffusion timestep $t$ in the diffusion schedule, the action-expert takes as input both the KV-cache from the sequence $\bm{o}_{\text{image}},\bm{o}_{\text{egomotion}},\textsc{Reason}]$ in the VLM and the embedded representation of the noisy control $\bm{a}_{t}$ (with the diffusion time $t$ also embedded and added to the feature). The expert then predicts the vector field $\mathbf{v}_{\Theta}(\bm{a}_{t},\bm{o},\textsc{Reason})$ by projecting the final layer feature through an MLP head, where $\Theta$ denotes the learnable parameters. We train the action-expert using a vanilla conditional flow matching loss (Lipman et al., 2023),
@@ 5.3.2 Reward Model
Grading Reasoning with Large Reasoning Models. To mitigate the issue where reasoning traces can exhibit hallucinations that produce plausible yet unsafe or causally inconsistent plans, we employ large reasoning models (LRMs) as automatic evaluators to provide scalable, high-quality feedback on reasoning quality. Inspired by recent advances in LLM alignment, where expert models serve as judges to provide scalable feedback (Bai et al., 2022; Lee et al., 2023a), we leverage state-of-the-art LRMs (e.g., DeepSeek-R1 (DeepSeek-AI, 2025), Cosmos-Reason (NVIDIA et al., 2025a)) as reasoning critics to evaluate the quality of reasoning traces generated by the VLA. We choose an LRM as the critic because, although such models may struggle to generate driving-specific reasoning due to limited embodiment priors, they exhibit strong verification and evaluation capabilities. In other words, even when generation in this domain is imperfect, their ability to assess logical soundness, causal alignment, and contextual consistency remains highly reliable (also known as the generation–verification gap (Song et al., 2024)). The resulting reward signal provides a continuous measure of reasoning quality, enabling RL to iteratively refine the model’s ability to generate grounded and logically consistent reasoning.
@@ 5.3.4 Post-Training Infrastructure
To conduct our RL experiments, we develop a customized version of the Cosmos-RL framework (NVIDIA, 2025) that is specifically designed for AV reasoning tasks. This system provides a scalable, modular infrastructure for large-scale multimodal RL and fits directly with other parts of the Alpamayo-R1 system. It supports distributed data loading, mixed-parallelism training, vLLM-based rollout generation (Kwon et al., 2023), and reward computation across multiple GPU nodes, enabling efficient, high-throughput policy optimization.
@@ 6 Experiments
In the following sections, we first present the evaluation protocol in Sec.˜6.1. Next, we illustrate how our reasoning-capable model contributes to an improved driving policy in Sec.˜6.2. In Sec.˜6.3 we further demonstrate the improvements in behavioral alignment achieved through RL. From Sec.˜6.4 to Sec.˜6.6 we conduct a comprehensive ablation study on the backbone model, the trajectory expert model, and the vision encoder to gain deeper insight into the effectiveness of our proposed methodology. Finally, we present an on-vehicle demonstration showcasing the real-world performance of our model.
@@ 6.1 Evaluation Protocol
Dataset. We train and evaluate models on internal driving data collected across diverse geographic regions in the US and EU, with all evaluation data strictly geo-fenced and held out from training regions to prevent information leakage. Our evaluation encompasses both nominal driving scenarios in dataset $\mathcal{D}_{\text{overall}}$ and challenging long-tail cases in $\mathcal{D}_{\text{hard}}$ to thoroughly test the model’s ability to handle rare, safety-critical events. In detail, the full training and evaluation dataset comprises 80,000 hours of driving data collected from multiple ego-vehicles operating in more than 1,700 cities in 25 countries. It encompasses diverse driving scenarios, including highway and urban environments, under various weather conditions, times of day, and traffic densities. The raw sensory inputs consist of video recordings from a surround-view seven-camera setup, accompanied by precise camera calibration parameters and ego-motion data. In this work, we focus on using two front-facing cameras as input: a front wide-angle camera with 120 field of view and a front telephoto camera with 30 field of view, providing complementary perspectives for both near-field and far-field scene understanding.
@@
Open-Loop Evaluation. For open-loop trajectory prediction, we evaluate models over a prediction horizon of 6 seconds, corresponding to the ego-vehicle’s planned waypoints. We use minADE and ADE as the evaluation metric. minADE is computed over 6 samples ($\text{minADE}_{6}$) and is defined as the minimum distance between the ground-truth future trajectory and the best-matching trajectory among 6 predictions generated by the model. ADE (Average Displacement Error) is the average distance between the predicted trajectory and the ground-truth trajectory across all future timesteps.
@@
    close encounter rate: percentage of scenarios where the ego vehicle experiences a close encounter with any other traffic agent;
@@
    AlpaSim score: average distance driven in km between events, where events correspond to offroad or close encounter occurrences;
  4. (4)
    AlpaSim score at fault: same as AlpaSim score but considering only close encounters where the ego vehicle is deemed responsible, i.e., excluding rear-end close encounters.
@@ 6.2 Policy Improvements from Reasoning
Open-Loop Improvements. As shown in Tab.˜6 (nominal scenarios) and Tab.˜7 (challenging scenarios), incorporating CoC reasoning yields substantial improvements in open-loop trajectory prediction in both settings. Without route information, AR1 achieves a minADE6 of 0.955m at 6s, a 4.1% improvement over the base model and outperforming both trajectory-only (0.971m) and meta-action (0.988m) baselines. With route information, the gains are more pronounced: AR1 achieves 0.794m, representing 4.8% improvement over the trajectory-only baseline (0.834m). Scaling to 3B parameters further improves performance, with AR1-3B achieving 0.908m (without route), demonstrating the benefits of increased model capacity for complex reasoning tasks. In challenging scenarios, the improvements are even larger, with AR1 achieving 0.868m, a 12% improvement over the trajectory-only baseline (0.994m).
@@
Closed-Loop Improvements. As shown in Tab.˜8, AR1 achieves a 35% reduction in off-road rate (11% vs 17%) and 25% reduction in close encounter rate (3% vs 4%) compared to the trajectory-only baseline. The overall AlpaSim score improves from 0.38 to 0.50, demonstrating that reasoning-based decision making improves safety in dynamic closed-loop scenarios. Fig.˜9 presents two qualitative examples that demonstrate that our model can successfully perform closed-loop driving in challenging scenarios within AlpaSim.
@@ 6.3 Improvements of Reasoning, Consistency, and Safety via RL Post-Training
Model Off-Road Rate $\downarrow$ (%) Close Encounter Rate $\downarrow$ (%) AlpaSim Score $\uparrow$ AlpaSim Score (at fault) $\uparrow$
@@ 6.4.1 Model Size Ablation
To investigate the impact of model capacity on driving performance, we first conduct baseline scaling experiments using general-purpose VLMs. Specifically, we evaluate three variants of our architecture with different backbone sizes: 0.5B, 3B, and 7B parameters. The 0.5B model uses a DINOv2 (Oquab et al., 2023) vision encoder combined with the Qwen2.5-0.5B (Qwen Team, 2024) language model, while the 3B and 7B models leverage Qwen2.5-VL-3B (Bai et al., 2025) and Qwen2.5-VL-7B (Bai et al., 2025), respectively. For this ablation study, all variants are trained on identical data with a reduced training budget compared to our main models, and evaluated on $\mathcal{D}_{\text{overall}}$ held-out test set without route information, using the minADE6 metric over a 6 s horizon.
As shown in Fig.˜12, we observe consistent improvements in open-loop performance as model size increases. The 7B model achieves a reduction of 11% in minADE6 compared to the baseline of 0.5B, demonstrating that scaling the vision-language backbone enables better scene understanding and trajectory prediction. While these results confirm the importance of model capacity, they are based on general-purpose VLMs without domain-specific pre-training. As we demonstrate in Sec.˜6.4.3, incorporating Physical AI-focused pre-training (via Cosmos-Reason, Sec.˜6.4.3) yields substantial further improvements, which is why our final Alpamayo-R1 models adopt Cosmos-Reason as their backbone.

### v2

@@ Abstract
End-to-end architectures trained via imitation learning have advanced autonomous driving by scaling model size and data, yet performance remains brittle in safety-critical long-tail scenarios where supervision is sparse and causal understanding is limited. We introduce Alpamayo-R1 (AR1), a vision–language–action model (VLA) that integrates Chain of Causation reasoning with trajectory planning for complex driving scenarios. Our approach features three key innovations: (1) the Chain of Causation (CoC) dataset, built through a hybrid auto-labeling and human-in-the-loop pipeline producing decision-grounded, causally linked reasoning traces aligned with driving behaviors; (2) a modular VLA architecture combining Cosmos-Reason, a vision-language model pre-trained for Physical AI, with a diffusion-based trajectory decoder that generates dynamically feasible trajectories in real time; (3) a multi-stage training strategy using supervised fine-tuning to elicit reasoning and reinforcement learning (RL) to enforce reasoning-action consistency and optimize reasoning quality. AR1 achieves up to a 12% improvement in planning accuracy on challenging cases compared to a trajectory-only baseline, with a 35% reduction in close encounter rate in closed-loop simulation. RL post-training improves reasoning quality by 45% and reasoning-action consistency by 37%. Model scaling from 0.5B to 7B parameters shows consistent improvements. On-vehicle road tests confirm real-time performance (99 ms latency) and successful urban deployment. By bridging interpretable reasoning with precise control, AR1 demonstrates a practical path towards Level 4 autonomous driving. Model weights are available at https://huggingface.co/nvidia/Alpamayo-R1-10B with inference code at https://github.com/NVlabs/alpamayo.
@@ 1 Introduction
The evolution of autonomous driving systems has witnessed a paradigm shift from traditional modular architectures (Urmson et al., 2008; Paden et al., 2016; Fan et al., 2018; Lefèvre et al., 2014) to end-to-end (E2E) driving frameworks (Bojarski et al., 2016; Hu et al., 2023; Jiang et al., 2023a; Weng et al., 2024; Wu, 2025), a transition increasingly embraced by industry. In contrast to modular designs that explicitly separate perception, prediction, and planning with hand-crafted intermediate representations, E2E approaches map raw sensor inputs directly to vehicle motion through jointly trained neural networks. This unified formulation eliminates manually engineered interfaces, enabling joint optimization and data-driven policy learning at scale. Recent advances in transformer-based architectures, coupled with large-scale driving datasets have further improved the overall performance and generalization of the E2E driving paradigm. Despite these successes, current E2E approaches remain fragile in handling long-tail and safety-critical situations, where sparse supervision and the need for high-level reasoning pose significant challenges.. Consequently, a significant gap persists between the capabilities of existing E2E models and the requirements for achieving robust Level-4 autonomy with driving-specific reasoning capabilities.
@@ 3 Building a Reasoning VLA Architecture
with each component conditioned on all previous ones. By default, the model is trained to predict the entire 6.4s-long future trajectory sequence
@@ 3.2.1 Vision Encoding
As an example, for $S_{x}=S_{y}=96,S_{z}=48$, and $p_{x}=p_{y}=p_{z}=8$, only 288 tokens are needed to represent one timestep of observations, irrespective of the number of cameras or their resolution. For a 7-camera vehicle setup, this equates to approximately $41.1$ tokens per image ($3.9\times$ less than single-image tokenization). Further, as we will show in Sec.˜6.7, this efficiency is achieved without major compromises to end-to-end driving metrics.
@@
Accordingly, AR1 also supports using multi-camera video tokenizers that directly encode entire sequences of camera observations from multiple timesteps. One example is Flex (Yang et al., 2025), which compresses a set of image tokens from multiple cameras and timesteps via full self-attention layers and a fixed set of query vectors, providing an explicit mechanism to control the magnitude of the information bottleneck. As will be shown in Sec.˜6.7, this approach can achieve an up to $20\times$ token compression rate (compared to single-image tokenization) while maintaining or even improving downstream driving metrics.
@@ 5.1 Action Modality Injection
Similar to $\pi_{0.5}$-KI (Driess et al., 2025), we adopt a separate action-expert to decode actions via flow matching (Janner et al., 2022; Lipman et al., 2023; Zhong et al., 2023; Jiang et al., 2023b). The action-expert follows the same Transformer architecture as the VLM, using the same number of attention heads and attention dimensions, but with a smaller hidden embedding size and MLP dimension for efficiency. At each diffusion timestep $t$ in the diffusion schedule, the action-expert takes as input both the KV-cache from the sequence $\bm{o}_{\text{image}},\bm{o}_{\text{egomotion}},\textsc{Reason}]$ in the VLM and the embedded representation of the noisy control $\bm{a}_{t}$ (with the diffusion time $t$ also embedded and added to the feature). The expert then predicts the vector field $\mathbf{v}_{\Theta}(\bm{a}_{t},\bm{o},\textsc{Reason})$ by projecting the final layer feature through an MLP head, where $\Theta$ denotes the learnable parameters. We train the action-expert using a vanilla conditional flow matching loss (Lipman et al., 2023),
@@ 5.3.2 Reward Model
Grading Reasoning with Large Reasoning Models. To mitigate the issue where reasoning traces can exhibit hallucinations that produce plausible yet unsafe or causally inconsistent plans, we employ large reasoning models (LRMs) as automatic evaluators to provide scalable, high-quality feedback on reasoning quality. Inspired by recent advances in LLM alignment, where expert models serve as judges to provide scalable feedback (Bai et al., 2022; Lee et al., 2023a), we leverage state-of-the-art LRMs (e.g., DeepSeek-R1 (DeepSeek-AI, 2025), Cosmos-Reason (NVIDIA et al., 2025a)) as reasoning critics to evaluate the quality of reasoning traces generated by the VLA. We choose an LRM as the critic because, although such models may struggle to generate driving-specific reasoning due to limited embodiment priors, they exhibit strong verification and evaluation capabilities. In other words, even when generation in this domain is imperfect, their ability to assess logical soundness, causal alignment, and contextual consistency remains highly reliable (also known as the generation–verification gap (Song et al., 2025)). The resulting reward signal provides a continuous measure of reasoning quality, enabling RL to iteratively refine the model’s ability to generate grounded and logically consistent reasoning.
@@ 5.3.4 Post-Training Infrastructure
To conduct our RL experiments, we develop a customized version of the Cosmos-RL framework (NVIDIA, 2025a) that is specifically designed for AV reasoning tasks. This system provides a scalable, modular infrastructure for large-scale multimodal RL and fits directly with other parts of the Alpamayo-R1 system. It supports distributed data loading, mixed-parallelism training, vLLM-based rollout generation (Kwon et al., 2023), and reward computation across multiple GPU nodes, enabling efficient, high-throughput policy optimization.
@@ 6 Experiments
In the following sections, we first present the evaluation protocol in Sec.˜6.1. Next, we illustrate how our reasoning-capable model contributes to an improved driving policy in Sec.˜6.2. In Sec.˜6.3 we further demonstrate the improvements in behavioral alignment achieved through RL. From Sec.˜6.5 to Sec.˜6.7 we conduct a comprehensive ablation study on the backbone model, the trajectory expert model, and the vision encoder to gain deeper insight into the effectiveness of our proposed methodology. Finally, we present an on-vehicle demonstration showcasing the real-world performance of our model.
@@ 6.1 Evaluation Protocol
Dataset. We train and evaluate models on internal driving data collected across diverse geographic regions in the US and EU, with all evaluation data strictly geo-fenced and held out from training regions to prevent information leakage. Our evaluation encompasses both nominal driving scenarios in dataset $\mathcal{D}_{\text{overall}}$ and challenging long-tail cases in $\mathcal{D}_{\text{hard}}$ to thoroughly test the model’s ability to handle rare, safety-critical events. In detail, the full training and evaluation dataset comprises 80,000 hours of driving data collected from multiple ego-vehicles operating in more than 2,500 cities in 25 countries. It encompasses diverse driving scenarios, including highway and urban environments, under various weather conditions, times of day, and traffic densities. The raw sensory inputs consist of video recordings from a surround-view seven-camera setup, accompanied by precise camera calibration parameters and ego-motion data. In this work, we focus on using two front-facing cameras as input: a front wide-angle camera with 120 field of view and a front telephoto camera with 30 field of view, providing complementary perspectives for both near-field and far-field scene understanding.
@@
Open-Loop Evaluation. For open-loop trajectory prediction, we evaluate models over a prediction horizon of 6.4 seconds, corresponding to the ego-vehicle’s planned waypoints. We use minADE and ADE as the evaluation metric. minADE is computed over 6 samples ($\text{minADE}_{6}$) and is defined as the minimum distance between the ground-truth future trajectory and the best-matching trajectory among 6 predictions generated by the model. ADE (Average Displacement Error) is the average distance between the predicted trajectory and the ground-truth trajectory across all future timesteps.
@@
    close encounter rate (all): percentage of scenarios where the ego vehicle experiences a close encounter with any other traffic agent;
  2. (2)
    close encounter rate (at-fault): same as close encounter rate but considering only close encounters where the ego vehicle is deemed responsible, i.e., excluding rear-end close encounters.
  3. (3)
@@
    AlpaSim score (all): average distance driven in km between events, where events correspond to offroad or close encounter occurrences;
@@
    AlpaSim score (at-fault): same as AlpaSim score but considering only close encounters where the ego vehicle is deemed responsible, i.e., excluding rear-end close encounters.
@@ 6.2 Policy Improvements from Reasoning
Open-Loop Improvements. As shown in Tab.˜6 (nominal scenarios) and Tab.˜7 (challenging scenarios), incorporating CoC reasoning yields substantial improvements in open-loop trajectory prediction in both settings. Without route information, AR1 achieves a minADE6 of 0.955m at 6.4s, a 4.1% improvement over the base model and outperforming both trajectory-only (0.971m) and meta-action (0.988m) baselines. With route information, the gains are more pronounced: AR1 achieves 0.794m, representing 4.8% improvement over the trajectory-only baseline (0.834m). Scaling to 3B parameters further improves performance, with AR1-3B achieving 0.908m (without route), demonstrating the benefits of increased model capacity for complex reasoning tasks. In challenging scenarios, the improvements are even larger, with AR1 achieving 0.868m, a 12% improvement over the trajectory-only baseline (0.994m).
@@
Closed-Loop Improvements. As shown in Tab.˜8, AR1 achieves a 35% reduction in close encounter rate (11% vs 17%) compared to the trajectory-only baseline, and a comparable off-road rate (4% vs 3%). The overall AlpaSim score improves from 0.38 to 0.50, demonstrating that reasoning-based decision making improves safety in dynamic closed-loop scenarios. Fig.˜9 presents two qualitative examples that demonstrate that our model can successfully perform closed-loop driving in challenging scenarios within AlpaSim.
@@ 6.3 Improvements of Reasoning, Consistency, and Safety via RL Post-Training
Close Encounter Rate Close Encounter Rate Off-Road Rate AlpaSim Score AlpaSim Score
Model all (%) $\downarrow$ at-fault (%) $\downarrow$ (%) $\downarrow$ all $\uparrow$ at-fault $\uparrow$
@@
6.4 Public Benchmark Evaluation
To enable reproducible evaluation and community comparison, we evaluate the publicly released Alpamayo-R1 model on the PhysicalAI-AV dataset (NVIDIA, 2025b) and the AlpaSim public scenario set (NVIDIA et al., 2025b). Alpamayo-R1-10B leverages Cosmos-Reason (NVIDIA et al., 2025a) as the VLM backbone with a 2B parameter diffusion-based trajectory decoder, while Alpamayo-R1-0.5B uses a smaller backbone for comparison. Both models generate Chain-of-Causation reasoning traces alongside trajectory predictions.
Table 10: Evaluation of Alpamayo-R1 on the public PhysicalAI-AV dataset (NVIDIA, 2025b). Open-loop results are evaluated on 644 examples from the PhysicalAI-AV evaluation set. Closed-loop results are evaluated on 920 scenarios from the PhysicalAI-AV NuRec dataset (NVIDIA, 2025c) using AlpaSim (NVIDIA et al., 2025b). All models predict CoC reasoning traces and trajectories. Closed-loop metrics are at-fault.
Model Open-Loop Closed-Loop (AlpaSim)
[…]
Open-Loop Results. We evaluate both models on 644 held-out examples from the PhysicalAI-AV dataset. As shown in Tab.˜10, Alpamayo-R1-10B achieves a minADE6 of 0.849m at 6.4s, a 7.0% improvement over Alpamayo-R1-0.5B (0.913m), demonstrating the benefits of scaling the VLM backbone.
[…]
@@ 6.5.1 Model Size Ablation
To investigate the impact of model capacity on driving performance, we first conduct baseline scaling experiments using general-purpose VLMs. Specifically, we evaluate three variants of our architecture with different backbone sizes: 0.5B, 3B, and 7B parameters. The 0.5B model uses a DINOv2 (Oquab et al., 2023) vision encoder combined with the Qwen2.5-0.5B (Qwen Team, 2024) language model, while the 3B and 7B models leverage Qwen2.5-VL-3B (Bai et al., 2025) and Qwen2.5-VL-7B (Bai et al., 2025), respectively. For this ablation study, all variants are trained on identical data with a reduced training budget compared to our main models, and evaluated on $\mathcal{D}_{\text{overall}}$ held-out test set without route information, using the minADE6 metric over a 6.4s horizon.
As shown in Fig.˜12, we observe consistent improvements in open-loop performance as model size increases. The 7B model achieves a reduction of 11% in minADE6 compared to the baseline of 0.5B, demonstrating that scaling the vision-language backbone enables better scene understanding and trajectory prediction. While these results confirm the importance of model capacity, they are based on general-purpose VLMs without domain-specific pre-training. As we demonstrate in Sec.˜6.5.3, incorporating Physical AI-focused pre-training (via Cosmos-Reason, Sec.˜6.5.3) yields substantial further improvements, which is why our final Alpamayo-R1 models adopt Cosmos-Reason as their backbone.